# Double Robustness of Local Projections

### v1

@@ 3.1 Model and assumptions
We consider a multivariate, stationary structural VARMA($1,\infty$) model that is local to an SVAR(1) model:
$y_{t}=Ay_{t-1}+H[I+T^{-\zeta}\alpha(L)]\varepsilon_{t}\quad\text{for all }t,$ (3.1)
where $y_{t}=(y_{1,t},\dots,y_{n,t})^{\prime}$ is $n$-dimensional, $\varepsilon_{t}=(\varepsilon_{1,t},\dots,\varepsilon_{m,t})^{\prime}$ is $m$-dimensional, $A$ is an $n\times n$ matrix, $H$ is an $n\times m$ matrix, and $\alpha(L)=\sum_{\ell=1}^{\infty}\alpha_{\ell}L^{\ell}$ is an $m\times m$ lag polynomial. We allow the number of shocks $m$ to potentially exceed the number of variables $n$, and vice versa. We will show below that equation (3.1) encompasses local-to-SVAR models with $p>1$ lags by writing them in companion form.
@@
    $\varepsilon_{t}\overset{i.i.d.}{\sim}(0_{m\times 1},D)$, where $D\equiv\operatorname*{diag}(\sigma_{1}^{2},\dots,\sigma_{m}^{2})$. For all $j=1,\dots,m$, $\sigma_{j}^{2}>0$ and $E(\varepsilon_{j,t}^{4})<\infty$.
@@
The assumption of shock homoskedasticity is made for analytical convenience, though we expect that our qualitative conclusions about the robustness of LP and the asymptotic bias of VAR will go through under various forms of conditional heteroskedasticity. The assumptions on $H$ correspond to recursive (also known as Cholesky) identification of the shock of interest $\varepsilon_{j^{*},t}$, with a unit effect normalization $H_{j^{*},j^{*}}=1$. A special case is when the shock is directly observed, which corresponds to ordering it first (i.e., $j^{*}=1$). It is a minor extension to allow for identification via external instruments (also known as proxies), as this just requires ordering the instrument $y_{1,t}$ first and adding an additional measurement error term to the first equation (Stock and Watson, 2018; Plagborg-Møller and Wolf, 2021).
The impulse response of interest is now defined as
@@
where $e_{i,n}$ denotes the $n$-dimensional unit vector with a 1 in position $i$.
@@
In particular, we can allow the estimation lag length $p$ to exceed the true minimal lag length $p_{0}$ of the model by setting $\check{A}_{\ell}=0$ for $\ell>p_{0}$. This extension will prove useful when we consider what happens as the lag length of the estimated VAR is increased.
@@
The local-to-SVAR model (3.1) with MA misspecification covers several empirically relevant types of misspecification. While essentially all modern discrete-time, linearized DSGE macro models have VARMA representations, they usually cannot be represented exactly as finite-order VAR models (e.g., Giacomini, 2013). [^5] See also Section 5 for concrete examples. Even if the true DGP were a finite-order VAR, dynamic misspecification of the estimation model can give rise to MA terms, for example due to under-specification of the lag length or failing to control for some of the variables in the true VAR system. Relatedly, MA terms may arise from a failure of invertibility of the shocks (Alessi, Barigozzi, and Capasso, 2011). VARMA representations can also arise from temporal or cross-sectional aggregation of finite-order VAR models, including contamination by classical measurement error (e.g., Granger and Morris, 1976; Lütkepohl, 1984). In all of these cases, if the number of lags used for estimating the VAR is chosen to be sufficiently large, then the MA remainder will be small in magnitude, which is consistent with the spirit of our locally misspecified model (3.1).
In the most general terms, our framework can accommodate arbitrary additive local misspecification of the form
$y_{t}=Ay_{t-1}+H\varepsilon_{t}+T^{-\zeta}\upsilon_{t},$
where $\upsilon_{t}$ is an unobserved, stationary, non-deterministic process that is independent of $\{\varepsilon_{t+\ell}\}_{\ell\geq 0}$. The parameter of interest $\theta_{h,T}$ is defined as the coefficient in a population projection of $y_{i^{*},t+h}$ onto $\varepsilon_{j^{*},t}$. The Wold decomposition theorem implies that $\upsilon_{t}$ has an MA representation (in a potentially expanded vector of shocks that includes $\varepsilon_{t}$), ultimately yielding a model of the form (3.1). This argument shows that our framework allows for VAR misspecification in the form of omitted nonlinear terms or stationary time-varying parameters, as long as (i) such misspecification is small relative to the linear, time-invariant VAR($p$) component; and (ii) we define the impulse response parameter as indicated above.
@@ 3.2 Estimators
    $y_{i^{*},t+h}=\hat{\beta}_{h}y_{j^{*},t}+\hat{\omega}_{h}^{\prime}\underline{y}_{j^{*},t}+\hat{\gamma}_{h}^{\prime}y_{t-1}+\hat{\xi}_{h,t}.$
    Recall from the previous subsection that if we are estimating an SVAR($p$) specification in the data $\check{y}_{t}$, then the vector $y_{t-1}$ actually contains $p$ lags $\check{y}_{t-1},\dots,\check{y}_{t-p}$. [^6] Controlling for lags in the LP is key to the double robustness property even if we directly observe the shock, if we want to allow for small contamination of the observed shock proxy.
@@
    The VAR estimator is defined as the response of $y_{i^{*},t+h}$ with respect to the $j^{*}$-th recursive innovation, where the magnitude of the innovation is normalized such that $y_{j^{*},t}$ increases by one unit on impact. That is, the estimator equals
@@
    and $\hat{C}_{\bullet,j^{*}}$ is the $j^{*}$-th column of the lower triangular Cholesky factor $\hat{C}$ of the covariance matrix $\hat{\Sigma}\equiv\frac{1}{T}\sum_{t=1}^{T}\hat{u}_{t}\hat{u}_{t}^{\prime}=\hat{C}\hat{C}^{\prime}$ of the residuals $\hat{u}_{t}\equiv y_{t}-\hat{A}y_{t-1}$. Again, in the case of an SVAR($p$) specification, the above formulas operate on the companion form.
It is well known that the two estimators coincide at the impact horizon: $\hat{\beta}_{0}=\hat{\delta}_{0}$ (see LABEL:thm:impact in LABEL:app:proof_details).
3.3 Asymptotic coverage
As in the univariate model in Section 2, the asymptotic representation of the LP estimator is invariant to misspecification, provided that $\zeta>1/4$.
@@ 3.3 Asymptotic coverage
The intuition for the robustness of LP is a straight-forward extension of the arguments given in the univariate case in Section 2 and so will not be repeated in detail here. LP is doubly robust in the sense that it is consistent if we correctly specify either the lagged controls in the outcome equation or the controls in the shock identification equation $\varepsilon_{j^{*},t}=y_{j^{*},t}-E[y_{j^{*},t}\mid\underline{y}_{j^{*},t},y_{t-1},y_{t-2},\dots]$. Even if both parts are misspecified, the two specification errors only affect the limiting distribution of $\hat{\beta}_{h}$ through their product. Hence, small specification errors of order $T^{-\zeta}$ impart an even smaller bias of order $T^{-2\zeta}$, which is asymptotically negligible relative to the standard deviation under our assumption $\zeta>1/4$.
In contrast, the VAR estimator is fragile, again as in the univariate case.
@@
$\hat{\delta}_{h}-\theta_{h,T}$ $=T^{-\zeta}\operatorname{aBias}(\hat{\delta}_{h})$
$\quad+\operatorname*{trace}\left\{S^{-1}\Psi_{h}HT^{-1}\sum_{t=1}^{T}\varepsilon_{t}\tilde{y}_{t-1}^{\prime}\right\}+\frac{1}{\sigma_{j^{*}}^{2}}e_{i^{*},n}^{\prime}A^{h}T^{-1}\sum_{t=1}^{T}\xi_{0,t}\varepsilon_{j^{*},t}+o_{p}(T^{-1/2}+T^{-\zeta}),$
@@
The VAR estimator suffers from bias of order $T^{-\zeta}$, while the stochastic terms of order $T^{-1/2}$ are the same as they would be in a correctly specified SVAR($p$) model. The first stochastic term captures sampling uncertainty in the reduced-form impulse responses $\hat{A}^{h}$, while the second term captures uncertainty in the structural impact response vector $\hat{\nu}$.
To study the coverage of LP and VAR confidence intervals, we first derive the asymptotic variances of the estimators.
@@
Under Assumption 3.1, the asymptotic covariance matrix for the order-$T^{-1/2}$ stochastic terms in the representations of the LP and VAR estimators in Propositions 3.1 and 3.2 is given by
$\operatorname{aVar}\left(\begin{pmatrix}\hat{\beta}_{h}\\ \hat{\delta}_{h}\end{pmatrix}\right)=\begin{pmatrix}\operatorname{aVar}(\hat{\beta}_{h})&\operatorname{aVar}(\hat{\delta}_{h})\\ \operatorname{aVar}(\hat{\delta}_{h})&\operatorname{aVar}(\hat{\delta}_{h})\end{pmatrix},$
where
$\operatorname{aVar}(\hat{\beta}_{h})$ $\equiv\frac{1}{\sigma_{j^{*}}^{2}}\left(e_{i^{*},n}^{\prime}A^{h}\overline{H}_{j^{*}}\overline{D}_{j^{*}}\overline{H}_{j^{*}}^{\prime}(A^{\prime})^{h}e_{i^{*},n}+\sum_{\ell=1}^{h}e_{i^{*},n}^{\prime}A^{h-\ell}\Sigma(A^{\prime})^{h-\ell}e_{i^{*},n}\right),$
$\operatorname{aVar}(\hat{\delta}_{h})$ $\equiv\frac{1}{\sigma_{j^{*}}^{2}}e_{i^{*},n}^{\prime}A^{h}\overline{H}_{j^{*}}\overline{D}_{j^{*}}\overline{H}_{j^{*}}^{\prime}(A^{\prime})^{h}e_{i^{*},n}+\operatorname*{trace}(\Psi_{h}\Sigma\Psi_{h}^{\prime}S^{-1}),$
and $\overline{D}_{j^{*}}\equiv\operatorname*{diag}(\sigma_{j^{*}+1}^{2},\dots,\sigma_{m}^{2})$.
@@
See Section C.1. ∎
For later reference, note that the corollary implies that $\hat{\beta}_{h}-\hat{\delta}_{h}$ is asymptotically independent of $\hat{\delta}_{h}$. This also follows from the general arguments of Hausman (1978), since Propositions 3.1 and 3.2 imply that the asymptotic covariance matrix of the LP and VAR estimators does not depend on the misspecification $\alpha(L)$, and the VAR estimator is the quasi-MLE (and thus efficient) in the correctly specified model. In particular, $\operatorname{aVar}(\hat{\delta}_{h})\leq\operatorname{aVar}(\hat{\beta}_{h})$.
The robustness of the LP confidence interval and fragility of the VAR confidence interval discussed in Section 2 carry over to the general model. We define the LP and VAR confidence intervals using the standard formula (2.5) in Section 2, but of course substituting the general expressions for the estimators and asymptotic variances defined in this section.
Corollary 3.2.
Under Assumption 3.1, $\lim_{T\to\infty}P(\theta_{h,T}\in\operatorname{CI}(\hat{\beta}_{h}))=1-a$. If, moreover, $\operatorname{aVar}(\hat{\delta}_{h})>0$ and $\operatorname{aBias}(\hat{\delta}_{h})\neq 0$, then $\lim_{T\to\infty}P(\theta_{h,T}\in\operatorname{CI}(\hat{\delta}_{h}))=\lim_{T\to\infty}\{1-r\left(T^{1/2-\zeta}b_{h};z_{1-a/2}\right)\}$, where $b_{h}\equiv\operatorname{aBias}(\hat{\delta}_{h})/\sqrt{\operatorname{aVar}(\hat{\delta}_{h})}$ and $r(b;c)\equiv P_{Z\sim N(0,1)}(|Z+b|>c)=\Phi(-c-b)+\Phi(-c+b)$.

### v2

@@ 2.1 Model and assumptions
Extending the forecasting model of Schorfheide (2005), we consider a multivariate, stationary structural VARMA($1,\infty$) model that is local to an SVAR(1) model:
$y_{t}=Ay_{t-1}+H[I+T^{-\zeta}\alpha(L)]\varepsilon_{t},\quad\text{for all }t,$ (2.1)
where the data vector $y_{t}=(y_{1,t},\dots,y_{n,t})^{\prime}$ is $n$-dimensional, the shock vector $\varepsilon_{t}=(\varepsilon_{1,t},\dots,\varepsilon_{m,t})^{\prime}$ is $m$-dimensional, $A$ is an $n\times n$ matrix, $H$ is an $n\times m$ matrix, $\alpha(L)=\sum_{\ell=1}^{\infty}\alpha_{\ell}L^{\ell}$ is an $m\times m$ lag polynomial, and $T$ denotes the sample size. We allow the number of shocks $m$ to potentially exceed the number of variables $n$, and vice versa. We show below that equation (2.1) encompasses local-to-SVAR models with $p>1$ lags by writing them in companion form.
The model (2.1) captures the idea that the time series dynamics of the data are well approximated by an autoregressive model driven by unobserved white noise shocks $\varepsilon_{t}$, but with a small amount of misspecification in the form of an MA process $T^{-\zeta}\alpha(L)\varepsilon_{t}$. The misspecification is asymptotically small in the sense that the MA coefficients converge to zero at the rate $T^{-\zeta}$, though the misspecification may still affect the properties of estimators, as shown by Schorfheide (2005) and as demonstrated below. We argue below that MA misspecification of this form can capture many empirically relevant types of dynamic misspecification. We consider local rather than global misspecification in the spirit of local power analysis (e.g., Rothenberg, 1984), since this makes the bias-variance trade-off between the VAR and LP estimators matter even asymptotically as the sample size $T$ diverges, allowing us to make tractable analytical comparisons between these two procedures.
@@
    $\varepsilon_{t}\overset{i.i.d.}{\sim}(0_{m\times 1},D)$, where $D\equiv\operatorname*{diag}(\sigma_{1}^{2},\dots,\sigma_{m}^{2})$, and the elements of $\varepsilon_{t}$ are mutually independent. For all $j=1,\dots,m$, $\sigma_{j}^{2}>0$ and $E(\varepsilon_{j,t}^{4})<\infty$.
@@
The assumption of shock homoskedasticity is made for analytical convenience, though we expect that our qualitative conclusions about the robustness of LP and the asymptotic bias of VAR will go through under various forms of conditional heteroskedasticity. The assumptions on $H$ correspond to recursive (also known as Cholesky) identification of the shock of interest $\varepsilon_{j^{*},t}$, with a unit effect normalization $H_{j^{*},j^{*}}=1$. A special case is when the shock is directly observed, which corresponds to ordering it first (i.e., $j^{*}=1$). It is a minor extension to allow for identification via external instruments, also known as proxies (Stock and Watson, 2018). Absolute summability of $\alpha(L)$ is a weak regularity condition ensuring the vector MA($\infty$) process $\alpha(L)\varepsilon_{t}$ is well-defined (Brockwell and Davis, 1991, Proposition 3.1.1). The significance of the assumption that misspecification vanishes faster than $T^{-1/4}$ will become clear below.
The impulse response of interest is defined as
@@
where $e_{i,n}$ denotes the $n$-dimensional unit vector with a 1 in position $i$. The first term in the parenthesis is the usual VAR impulse response formula, while the second term arises from the MA component. Importantly, and consistent with our focus on the consequences of dynamic misspecification, we do not treat the VAR misspecification as non-classical measurement error that should be ignored for structural analysis; instead, the true causal model has a VARMA form (with small but potentially non-zero MA terms), and we care about the full transmission mechanism of shocks in this model.
@@
In particular, we can allow the estimation lag length $p$ to exceed the true minimal lag length $p_{0}$ of the model by setting $\check{A}_{\ell}=0$ for $\ell>p_{0}$. This fact will prove useful when we consider what happens as the lag length of the estimated VAR is increased.
@@
Our local-to-SVAR model (2.1) with MA misspecification covers several empirically relevant types of model misspecification. While essentially all modern discrete-time, linearized DSGE macro models have VARMA representations, they usually cannot be represented exactly as finite-order VAR models (e.g., Kilian and Lütkepohl, 2017, Chapter 6.2). Even if the true DGP were a finite-order VAR, dynamic misspecification of the estimation model can give rise to MA terms, for example due to under-specification of the lag length or failing to control for some of the variables in the true system. Relatedly, MA terms may appear because of a failure of invertibility of the shocks (Alessi, Barigozzi, and Capasso, 2011). VARMA representations can also arise from temporal or cross-sectional aggregation of finite-order VAR models, including contamination by classical measurement error (Granger and Morris, 1976; Lütkepohl, 1984). In all of these cases, if the number of lags used for estimating the VAR is chosen to be sufficiently large, then the MA remainder will be small, consistent with the spirit of our locally misspecified model (2.1).
In terms of structural shock identification, our framework accommodates both the case of a well-identified shock (but misspecification in other parts of the model) and misspecification in the shock identification itself. Key to this generality is that we allow the $m\times m$ MA polynomial $\alpha(L)$ to be arbitrary. To see this, consider the case $j^{*}=1$, so interest centers on the dynamic causal effects of the first shock $\varepsilon_{1,t}$. If the first row of $\alpha(L)$ is zero, then $\varepsilon_{1,t}$ is well-identified as the reduced-form residual in the first equation of the VAR. If the first row of $\alpha(L)$ is non-zero, then the reduced-form residual will be contaminated by lagged shocks, thus allowing for the possibility that shock identification is not entirely accurate.
We conjecture, but do not prove, that our framework can also be extended to accommodate omitted nonlinearities and time-varying parameters, as long as those features are small compared to the linear, time-invariant model components. [^1] Specifically, we conjecture that our proofs can be extended to allow for arbitrary additive local misspecification of the form $y_{t}=Ay_{t-1}+H\varepsilon_{t}+T^{-\zeta}\upsilon_{t},$ where $\upsilon_{t}$ is an unobserved, stationary, non-deterministic process that is independent of $\{\varepsilon_{t+\ell}\}_{\ell\geq 0}$. Importantly, the parameter of interest $\theta_{h,T}$ must be defined as the coefficient in a population projection of $y_{i^{*},t+h}$ onto $\varepsilon_{j^{*},t}$. The Wold decomposition theorem implies that $\upsilon_{t}$ has an MA representation, ultimately yielding a model of the form (2.1). However, since the orthogonalized Wold innovations will not satisfy all the independence conditions in Assumption 2.1, our current proofs do not apply directly.
@@ 2.2 Estimators
    $y_{i^{*},t+h}=\hat{\beta}_{h}y_{j^{*},t}+\hat{\omega}_{h}^{\prime}\underline{y}_{j^{*},t}+\hat{\gamma}_{h}^{\prime}y_{t-1}+\hat{\xi}_{i^{*},h,t},$ (2.3)
    where $\hat{\xi}_{i^{*},h,t}$ is the least-squares residual. Recall from the previous subsection that if we are estimating an SVAR($p$) specification in the data $\check{y}_{t}$, then the vector $y_{t-1}$ actually contains $p$ lags $\check{y}_{t-1},\dots,\check{y}_{t-p}$.
@@
    The VAR estimator is defined as the response of $y_{i^{*},t+h}$ with respect to the $j^{*}$-th recursively orthogonalized innovation, where the magnitude of the innovation is normalized such that $y_{j^{*},t}$ increases by one unit on impact:
@@
    and $\hat{C}_{\bullet,j^{*}}$ is the $j^{*}$-th column of the lower triangular Cholesky factor $\hat{C}$ of the covariance matrix $\hat{\Sigma}\equiv\frac{1}{T}\sum_{t=1}^{T}\hat{u}_{t}\hat{u}_{t}^{\prime}=\hat{C}\hat{C}^{\prime}$ of the residuals $\hat{u}_{t}\equiv y_{t}-\hat{A}y_{t-1}$. Again, in the case of an SVAR($p$) specification, the above formulae operate on the companion form.
Note that the two estimators coincide at the impact horizon: $\hat{\beta}_{0}=\hat{\delta}_{0}$ (see LABEL:thm:impact in LABEL:app:proof_details).
It is well-known that conventional confidence intervals based on both these estimators would have correct asymptotic coverage in a well-specified VAR model. However, the presence of the additional MA term in the model (2.1) means that, in principle, both the LP and VAR estimators ought to control for infinitely many lags of the data, rather than just one. Nevertheless, as we will now establish, this dynamic misspecification has much more serious consequences for the VAR procedure than for LP.
3 Robust local projections, fragile VARs
This section shows that the conventional LP confidence interval is robust to large amounts of misspecification. In contrast, the conventional VAR confidence interval has fragile coverage, except when it is asymptotically as wide as the LP interval, as will be the case with sufficiently large lag length.
3.1 Large-sample distributions and confidence interval coverage
We begin by characterizing the large-sample distributions of the LP and VAR estimators.
The robustness of LPs.
Our first main result establishes that the large-sample distribution of the LP estimator is invariant to misspecification.
@@ 3.1 Large-sample distributions and confidence interval coverage
In words, the asymptotic behavior of LP does not depend on the misspecification parameters $\alpha(L)$ and $\zeta$, provided $\zeta>1/4$ as imposed in Assumption 2.1. Though this robustness property of LP is with respect to local (i.e., asymptotically vanishing) misspecification, it is still quantitatively meaningful, given that MA terms of order $T^{-\zeta}$ with $\zeta\in(1/4,1/2)$ can be detected with probability 1 asymptotically by conventional VAR model specification tests, such as the Hausman test considered in Section 3.2.
Why is LP robust to misspecification of such large magnitude? We will offer two mathematically equivalent pieces of intuition, with our discussion throughout deliberately heuristic. The classic omitted variable bias (OVB) formula suggests that the bias of the LP impulse response estimator $\hat{\beta}_{h}$ in the regression (2.3) is proportional to the product of two factors: (i) the direct effect of omitted lags on $y_{i^{*},t+h}$, and (ii) the covariance of the residualized regressor of interest $y_{j^{*},t}-E[y_{j^{*},t}\mid\underline{y}_{j^{*},t},y_{t-1}]$ with the omitted lags. The factor (i) is of order $T^{-\zeta}$ in our local-to-SVAR model (2.1). The factor (ii) is also of order $T^{-\zeta}$, since the residualized regressor equals $\varepsilon_{j^{*},t}+O_{p}(T^{-\zeta})$ under Assumption 2.1(iii), and the shock $\varepsilon_{j^{*},t}$ is uncorrelated with any lagged data. Hence, the OVB is of order $T^{-2\zeta}=o(T^{-1/2})$ when $\zeta>1/4$, so the bias of the estimator is negligible relative to the standard deviation (which is of order $T^{-1/2}$, as in the correctly specified case). This argument relies on the LP regression controlling for the most important lags of the data (i.e., $y_{t-1}$); without lagged controls, one or both factors in the OVB formula may not be small (González-Casasús and Schorfheide, 2024).
The preceding intuition is a special case of the double robustness property of partially linear regressions, see Example 1.1 in Chernozhukov et al. (2018) and Example 1 in Chernozhukov et al. (2022). We will now argue that this property applies also to LP, again settling for a heuristic argument. For notational simplicity, set $j^{*}=1$ so $\underline{y}_{j^{*},t}=0$. Consider any dynamic model (for example a VARMA($p,q$)) that implies the following local projection representation:
[…]
$E[\{y_{i^{*},t+h}-\theta_{0,h}y_{1,t}-\gamma_{0}(y^{t-1})\}\{y_{1,t}-\nu_{0}(y^{t-1})\}]=0.$
If we evaluate the moment on the left-hand side at arbitrary functions $\gamma(\cdot)$ and $\nu(\cdot)$ rather than at the true ones $\gamma_{0}(\cdot)$ and $\nu_{0}(\cdot)$, a simple calculation shows that it equals $E[\{\gamma_{0}(y^{t-1})-\gamma(y^{t-1})\}\{\nu_{0}(y^{t-1})-\nu(y^{t-1})\}]$. [^2] We can write the moment as $E[\{y_{i^{*},t+h}-\theta_{0,h}y_{1,t}-\gamma_{0}(y^{t-1})+\gamma_{0}(y^{t-1})-\gamma(y^{t-1})\}\{y_{1,t}-\nu(y^{t-1})\}]=E[\{\gamma_{0}(y^{t-1})-\gamma(y^{t-1})\}\{y_{1,t}-\nu_{0}(y^{t-1})+\nu_{0}(y^{t-1})-\nu(y^{t-1})\}]$, since $y_{i^{*},t+h}-\theta_{0,h}y_{1,t}-\gamma_{0}(y^{t-1})=\xi_{i^{*},h,t}$ is independent of $y^{t}$ (orthogonality would suffice if $\nu(\cdot)$ were linear). The claim now follows from $E[y_{1,t}-\nu_{0}(y^{t-1})\mid y^{t-1}]=0$ by definition of $\nu_{0}(\cdot)$. Hence, the moment condition is satisfied at the true impulse response parameter $\theta_{0,h}$ as long as either $\gamma=\gamma_{0}$ or $\nu=\nu_{0}$, making the LP estimator doubly robust: it is consistent if we correctly specify either the controls $\gamma(y^{t-1})$ in the outcome equation or the controls $\nu(y^{t-1})$ in the implicit first-stage regression that isolates the shock $\varepsilon_{j^{*},t}=y_{j^{*},t}-\nu(y^{t-1})$. Because of double robustness, and as argued more generally by Chernozhukov et al. (2018) (and confirmed by our proof), it turns out that estimation error in $\gamma_{0}$ and $\nu_{0}$ only affects the asymptotic distribution of $\hat{\beta}_{h}$ through the product of the estimation errors $\|\hat{\gamma}-\gamma_{0}\|\times\|\hat{\nu}-\nu_{0}\|$. In our local-to-SVAR model (2.1), both terms in this product are of order $T^{-\zeta}$ due to the omitted lags. The product is then of order $T^{-2\zeta}=o(T^{-1/2})$ and thus asymptotically negligible, consistent with our earlier intuition.
The fragility of VARs.
In contrast to LP, the VAR estimator is fragile, as it is subject to non-negligible asymptotic bias under misspecification.
@@
$\hat{\delta}_{h}-\theta_{h,T}$ $=\operatorname*{trace}\left\{S^{-1}\Psi_{h}HT^{-1}\sum_{t=1}^{T}\varepsilon_{t}\tilde{y}_{t-1}^{\prime}\right\}+\frac{1}{\sigma_{j^{*}}^{2}}e_{i^{*},n}^{\prime}A^{h}T^{-1}\sum_{t=1}^{T}\xi_{0,t}\varepsilon_{j^{*},t}$
$\quad+T^{-\zeta}\operatorname{aBias}(\hat{\delta}_{h})+o_{p}(T^{-1/2}+T^{-\zeta}),$
@@
The convergence rate $T^{-\min\{1/2,\zeta\}}$ of the VAR estimator is potentially slower than the $T^{-1/2}$ rate achieved by LP. This is because the VAR estimator suffers from bias of order $T^{-\zeta}$, while the stochastic terms of order $T^{-1/2}$ are the same as they would be in a correctly specified SVAR($p$) model. [^3] The first stochastic term captures sampling uncertainty in the reduced-form impulse responses $\hat{A}^{h}$, while the second term captures uncertainty in the structural impact response vector $\hat{\nu}$. The VAR bias is only asymptotically negligible if $\zeta>1/2$, a much smaller degree of robustness than shown above for LP. The case $\zeta=1/2$ is of particular interest, as then the bias and standard deviation are of the same asymptotic order (see also Schorfheide, 2005). MA terms of order $T^{-1/2}$ can be detected with asymptotic probability strictly between 0 and 1 by specification tests, as will be shown in Section 3.2.
The asymptotic bias is due to two forces: first, the coefficient matrix $\hat{A}$ is biased due to the endogeneity caused by the MA terms, and second, the VAR estimator extrapolates the $h$ impulse response based on a parametric formula $\hat{A}^{h}$ that does not hold exactly in the true VARMA model (2.1). This is more easily seen in the special case of a univariate model $y_{t}=\rho y_{t-1}+[1+T^{-\zeta}\alpha(L)]\varepsilon_{t}$ with $n=m=1$, in which case
$\operatorname{aBias}(\hat{\delta}_{h})\equiv\underbrace{h\rho^{h-1}}_{\frac{\partial(\rho^{h})}{\partial\rho}}\underbrace{(1-\rho^{2})\sum_{\ell=1}^{\infty}\rho^{\ell-1}\alpha_{\ell}}_{\operatorname{aBias}(\hat{\rho})=\frac{\operatorname*{Cov}(\alpha(L)\varepsilon_{t},\tilde{y}_{t-1})}{\operatorname*{Var}(\tilde{y}_{t-1})}}-\underbrace{\sum_{\ell=1}^{h}\rho^{h-\ell}\alpha_{\ell}}_{\theta_{h,T}-\rho^{h}},$
where $\hat{\rho}=\hat{A}$ is the AR(1) coefficient from an OLS regression of $y_{t}$ on $y_{t-1}$. [^4] Lag augmentation of the VAR impulse response estimator as in Inoue and Kilian (2020) may reduce the first term in the bias formula, but it does not affect the second term.
Confidence intervals.
The preceding results imply that the conventional LP confidence interval is robust to misspecification while the conventional VAR interval is not. We define the level-($1-a$) LP and VAR confidence intervals using the standard formulae:
$\operatorname{CI}(\hat{\beta}_{h})\equiv\left[\hat{\beta}_{h}\pm z_{1-a/2}\sqrt{\operatorname{aVar}(\hat{\beta}_{h})/T}\right],\quad\operatorname{CI}(\hat{\delta}_{h})\equiv\left[\hat{\delta}_{h}\pm z_{1-a/2}\sqrt{\operatorname{aVar}(\hat{\delta}_{h})/T}\right].$ (3.1)
Here $z_{1-a/2}$ is the $21-a/$ 2 quantile of the standard normal distribution, and $\operatorname{aVar}(\hat{\beta}_{h})$ and $\operatorname{aVar}(\hat{\delta}_{h})$ are the asymptotic variances of the leading (order-$T^{-1/2}$) stochastic terms in the representations of the LP and VAR estimators in Propositions 3.1 and 3.2; explicit formulae for the asymptotic variances are given in Corollary A.2 in Section A.3, which also implies that $\operatorname{aVar}(\hat{\beta}_{h})\geq\operatorname{aVar}(\hat{\delta}_{h})$. None of the results below would change if we replaced the asymptotic variances with the conventional consistent estimates of these (that assume correct specification, as implemented in standard econometric software packages). [^5] Under Assumption 2.1, homoskedastic OLS standard errors suffice. In practice we recommend heteroskedasticity-robust standard errors or the bootstrap, as in Montiel Olea and Plagborg-Møller (2021).
@@
Under Assumption 2.1, $\lim_{T\to\infty}P(\theta_{h,T}\in\operatorname{CI}(\hat{\beta}_{h}))=1-a$. If moreover $\operatorname{aVar}(\hat{\delta}_{h})>0$ and $\operatorname{aBias}(\hat{\delta}_{h})\neq 0$, then $\lim_{T\to\infty}P(\theta_{h,T}\in\operatorname{CI}(\hat{\delta}_{h}))=\lim_{T\to\infty}\{1-r\left(T^{1/2-\zeta}b_{h};z_{1-a/2}\right)\}$, where $b_{h}\equiv\operatorname{aBias}(\hat{\delta}_{h})/\sqrt{\operatorname{aVar}(\hat{\delta}_{h})}$, $r(b;c)\equiv P_{Z\sim N(0,1)}(|Z+b|>c)=\Phi(-c-b)+\Phi(-c+b)$, and $\Phi(\cdot)$ is the standard normal distribution function.
@@
Considering separately the three cases $\zeta\in(1/4,1/2)$, $\zeta=1/2$, and $\zeta>1/2$, the result is an immediate consequence of Propositions 3.1 and 3.2. ∎
LP robustly controls coverage, while the VAR confidence interval generically has coverage converging to zero for $\zeta\in(1/4,1/2)$, and strictly below the nominal level $1-a$ for $\zeta=1/2$. Intuitively, the VAR confidence interval has the right width (the same as in the correctly specified case) but the wrong location due to the bias, thus causing coverage distortions.
3.2 Hausman misspecification test
To aid in interpreting the magnitude of the local misspecification in our set-up, we consider a Hausman (1978) test of correct specification of the VAR model that compares the VAR and LP impulse response estimates. This test rejects for large values of $\sqrt{T}|\hat{\beta}_{h}-\hat{\delta}_{h}|/\sqrt{\operatorname{aVar}(\hat{\beta}_{h})-\operatorname{aVar}(\hat{\delta}_{h})}$. A test of this kind was proposed by Stock and Watson (2018) in the context of testing for invertibility.
Proposition 3.3.
where $b_{h}$ and $r(\cdot,\cdot)$ were defined in Corollary 3.1.